# Dark Dimension and the Effective Field Theory limit

### v1

@@ Abstract
with no need for fine-tuning. In [1] we showed that the right hand side of (5) also contains UV-sensitive terms (see the first line in the right hand side of Eq. (Dark Dimension and the Effective Field Theory limit) below), that were missed in previous literature. Therefore, the matching between the prediction (4) for the vacuum energy $\Lambda_{cc}$, that comes from swampland conjectures in string theory, and the corresponding result in the low energy field theory limit is not as straightforward as it was thought on the basis of the result (5). In this respect, it is worth to stress again (see the paragraph below Eq. (3)) that such a matching is crucial for the DD scenario.
@@
Finally, we would like to spend one word on the general framework in which in our opinion the DD scenario should be framed. In [1] and in the present work we considered the contribution to the vacuum energy that comes from the bulk. The reason is that in the paper that contains the DD proposal [2] as well as in the paper [3] that attempts to confute the results of our work [1] this is the only part of the vacuum energy taken into account. Naturally, there is also the contribution coming from the Standard Model, that lives on a 3-brane. Since the SM physical cutoff $\Lambda_{\rm SM}$ is at least of the TeV order, the latter cannot be ignored and viability of the DD scenario is subject to the existence of a physical mechanism that could dispose of this too large contribution. In our opinion, the suppression mechanism for both bulk and brane (SM and possibly BSM) contributions should be driven by a global renormalization group (RG) flow that takes into account the whole $(4+1)$D theory, including the SM in the 3-brane. The UV-complete theory (string theory/quantum gravity) should provide the boundary (UV) value of the vacuum energy (more generally the boundary values of all the physical parameters), and RG equations should dictate how the measured vacuum energy is reached in the IR [32]. Assessing the presence/absence of UV-sensitive terms in loop calculations is crucial to determine this UV-IR flow. Work is in progress in this direction [33].
@@ Acknowledgments
We thank G. Dvali and M. Scalisi for useful discussions. The work of CB has been supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567), and by the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR). The work of VB, FC and AP is carried out within the INFN project QGSKY.

### v2

@@ Abstract
with no need for fine-tuning. In [1] we showed that the right hand side of (5) also contains UV-sensitive terms (see the first line in the right hand side of Eq. (Dark Dimension and the Effective Field Theory limit) below), that were missed in previous literature [^3] The presence of similar UV-sensitive terms in the calculation of the effective potential of a scalar theory was first hinted in [24]. Moreover, it was shown in [25] that the presence of Fayet-Iliopulos terms induces further divergences.. Therefore, the matching between the prediction (4) for the vacuum energy $\Lambda_{cc}$, that comes from swampland conjectures in string theory, and the corresponding result in the low energy field theory limit is not as straightforward as it was thought on the basis of the result (5). In this respect, it is worth to stress again (see the paragraph below Eq. (3)) that such a matching is crucial for the DD scenario.
@@
Finally, we would like to spend one word on the general framework where, in our opinion, the DD scenario should be framed. In [1] and in the present work we considered the contribution to the vacuum energy that comes from the bulk. The reason is that in the paper that contains the DD proposal [2], as well as in the paper [3] that attempts to confute the results of our work [1], this is the only part of the vacuum energy taken into account. Naturally, there is also a contribution from the Standard Model, that, as already said, lives on a 3-brane. Since the SM physical cutoff $\Lambda_{\rm SM}$ is at least of the TeV order, this contribution cannot be ignored. In this respect, it is important to stress that it has been already objected in the past (well before the formulation of the DD proposal) that the SM contribution to the vacuum energy seems to invalidate any attempt to get a small cosmological constant in theories with compact extra dimensions [34, 35].
The viability of the DD scenario is certainly subject to the existence of a physical mechanism that could dispose of this too large contribution. We expect that, for both bulk and brane contributions, such a mechanism (if any) should be provided by the piling up of quantum fluctuations in the UV $\to$ IR renormalization group (RG) flow that takes into account the whole $(4+1)$D theory, including the SM in the 3-brane. In such a framework, the UV-complete theory (string theory/quantum gravity) should provide the boundary (i.e.  UV) value of the vacuum energy (more generally the boundary values of all the physical parameters), so that RG equations should dictate how the measured vacuum energy is reached in the IR (see [36] for examples where this mechanism is realized in $d=3$ and $d=4$ scalar theories). Work is in progress in this direction [37].
@@ Acknowledgments
We thank S. Abel, G. Dvali, H. P. Nilles and M. Scalisi for useful discussions. The work of CB has been partly supported by the Basic Science Research Program through the National Research Foundation of Korea (NRF) funded by the Ministry of Education, Science and Technology (NRF-2022R1A2C2003567), and partly by the European Union – Next Generation EU through the research grant number P2022Z4P4B “SOPHYA - Sustainable Optimised PHYsics Algorithms: fundamental physics to build an advanced society” under the program PRIN 2022 PNRR of the Italian Ministero dell’Università e Ricerca (MUR). The work of VB, FC and AP is carried out within the INFN project QGSKY.